Use the following information to answer the next few questions:
The production manager for the Coory soft drink company is considering the production of 2 kinds of soft drinks: regular (R) and diet (D). Two of her limited resources are production times (8 hours = 480 minutes per day) and syrup (1 of her ingredients) limited to 675 gallons per day. To produce a regular case requires 2 minutes and 5 gallons of syrup, which a diet case needs 4 minutes and 3 gallons of syrup. Profits for regular soft drink are $3 per case and profits for diet soft drink are $2 per case. Set this up as a linear programming model and solve.

The total profit for the above model's diet drink only alternative non-optimal solution is:

Question 5 options:

120

96

480

135

240
Answer #1

Cases of Let ex and eys be the number of and regular broduced diet soft drink to be As regular case requires 2 minutes while

We can use excel solvar to set up the linear programming problem as shown -

E12 : X A B с F 1 х Y Condition Value 2 3 4. Profit =3*B3 =B3 5 T =2*C3 =2* C3 =3*C3 =SUM(B4:04) =SUM(B5:C5) =SUM(B6:06) 240Then set up the solver as shown -

File Home Insert Draw Page Layout Formulas Data Review View Help Power Pivot Share LÈ From Text/CSV E Queries & Connections G

Note that you can access 'Solver' from 'Data' tab as shown above.

This should give you following output -

X = 90 and Y = 75.

Thus, maximum profit = Z = 3(90) + 2(75)

= $420

Now, if you want only diet alternative profit, so set X = 0, to set up the LP as  -

D4 X fr B с D E F 0 Solver Parameters Х X Y Condition Value А. 1 2 3 4 Profit 5 T 0 Set Objective: $D$4 4 0 0 oi 0 0 To: 0<%3

Now, this would give you the output -

X Y
0 120
Profit 0 240 240
T 0 240 240
S 0 360 360

So, you produce 120 cases of diet coke which gives you a maximum profit of -

Z = 240

Thus, last option is correct.

_____________________________